OpenStudy (anonymous):

Each system of DE is a model for 2 species that either compete for the same resources or cooperate for mutual benefit. \[\frac{dx}{dt}=0.12x-0.0006x^2+0.00001xy\]

OpenStudy (anonymous):

I think you are missing some info

OpenStudy (anonymous):

\[\frac{dy}{dt}=0.08x+0.00004xy\]

OpenStudy (anonymous):

I have to equations, one with dx/dt and the other with dy/dt. What's the main idea behind that?

OpenStudy (anonymous):

this would be totally okay for me if it were linear, but...

OpenStudy (anonymous):

It's only calc II

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Yes. I think I'm supposed to only decide whether each system describes a competition or cooperation.

OpenStudy (anonymous):

hm... this is seeming strange to me all this terminology is new, though I should know it if it's calc II

OpenStudy (anonymous):

The chapter is differential equations and the section is "predator-prey systems" \[\frac{dR}{dt}=kR\] k is a positive constant \[\frac{dW}{dt}=-rW\] r is a positive constant \[\frac{dR}{dt}=kR-aRW\] \[\frac{dW}{dt}=-rW=bRW\]

OpenStudy (anonymous):

the last 2 equation are the predaor-prey equations

OpenStudy (fwizbang):

Look at the xy term in both of the equations. If both coefficients are positive, then rate of change of x and y both increase, which means that each species benefits from the other's presence...... a cooperation. If both terms were negative, then each would increase its prospects by driving the other out of existence, which would be competition.